OpenStudy (anonymous):

find the median mode and range 34,17,10,23,21,15

OpenStudy (anonymous):

5th grade math

OpenStudy (anonymous):

median is the middle value between the organized set. mode is the most reoccuring number in the set range is the largest number in the set minus the smallest

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so can you help me solve it

OpenStudy (anonymous):

no

OpenStudy (anonymous):

ok then

OpenStudy (anonymous):

You can organize the set and you will be able to find the values.

OpenStudy (anonymous):

but i really dont know how to do it

OpenStudy (anonymous):

so i put then it the least to greatest

OpenStudy (anonymous):

right

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so the try to find which one is your middle value.

OpenStudy (anonymous):

this may be trick since you have an even number of numbers

OpenStudy (anonymous):

you will have to take the two middle values and add them and divide by 2

OpenStudy (anonymous):

Mean = 34+17+10+23+21+15 / 6 = 20 Reorder to get the median, it is the middle value usually but in this case you do 10, 15, 17, 21, 23, 34 so you take 17 and 21 and take average of these two like (17+21)/2 = 38/2 = 19 Mode = none because no value is repeating Range is the largest number minus the lowest number so, 34-10=24

OpenStudy (anonymous):

Mode-none Range-24 median-19